### v1

@@ 1 Introduction
Several prior studies have investigated estimators that target the cluster-ATE or individual-ATE as two marginal estimands in CRTs. Without baseline covariate adjustment, Table 3 in Kahan et al.4 provided a recipe for consistent estimators using either the individual-level data or cluster-summary data; they recommended the (weighted) independence generalized estimating equations (GEE) when the cluster size is informative. Wang et al.11 pointed out that the probability limit of the treatment coefficient estimator from GEE with an exchangeable working correlation structure is neither the cluster-ATE nor the individual-ATE. Similar findings also apply to parallel-arm CRTs with a baseline period.12 With covariate adjustment, Su and Ding 13 established the asymptotic theory for a class of linearly-adjusted analysis of covariance estimators for estimating the weighted average treatment effects in CRTs, assuming working independence. Balzer et al.14, 15 and Benitez et al.10 studied targeted maximum likelihood estimators with machine learning methods for covariate adjustment in CRTs, and extended their methods to address missing outcomes;16 A further extension allowing for covariates missing not at random is given in Wang et al.17 Wang et al.18 developed efficient covariate-adjusted estimators for estimating both the cluster-ATE and individual-ATE, allowing for informative within-cluster subsampling. Despite these recent developments, the most common practice for analyzing CRTs relies on fitting regression models suitable for clustered data, including the (generalized) linear mixed models and GEE. For example, Fiero et al.19 reviewed 86 CRTs published between August 2013 and July 2014 and found that 52% used mixed models and 16% used GEE in their primary analyses. In a review of 100 CRTs from the United Kingdom’s National Institute for Health Research online Journals Library (from January 1997 to July 2021), 80% of the studies adopted a linear mixed or generalized linear mixed model for data analysis.20 For practitioners, guidance has been limited on how to draw robust inference about the cluster-ATE and the individual-ATE using common regression models for clustered data, irrespective of choices regarding covariate adjustment, working correlation and random-effects specification.
@@ 2 Formalizing the treatment effect estimands
We consider a standard parallel-arm CRT, where individuals are nested within clusters. We assume that data observations are collected through $m$ clusters, each of which contains $N_{i}$ individuals that we include in a study. We assume the cluster size $N_{i}$ is the study population of interest in each cluster and no further within-cluster subsampling is considered; methods that address the generalization to potentially unobserved cluster population can be found in Wang et al.18 and may require the additional knowledge of the source population size in each cluster. We let $A_{i}\in\{0,1\}$ be the randomized cluster-level treatment indicator, with $A_{i}=1$ indicating the assignment to the treatment condition and $A_{i}=0$ to usual care. Under the potential outcomes framework, we define $\{Y_{ij}(1),Y_{ij}(0)\}$ as a pair of potential outcomes for each individual $j\in\{1,\dots,N_{i}\}$ under the treatment and usual care conditions, respectively. Writing $f(a,b)$ as a pre-specified contrast function, we define a general class of weighted average treatment effect in CRTs as
@@
Although estimand (3) does not appear to be an individual-average quantity from a first look, its finite-population analogue can be defined as $\sum_{i=1}^{m}\sum_{j=1}^{N_{i}}Y_{ij}(a)/\sum_{i=1}^{m}N_{i}$ (by replacing the population-level expectation with an empirical one)13 and clearly reveals its feature. The individual-ATE addresses the expected change in outcome associated with treatment across the population of individuals across clusters, and can be relevant when the intervention is targeting either the cluster or the individual (despite cluster randomization).
@@
    (Conceptual difference) The cluster-ATE and the individual-ATE estimands focus on different units of inference,21 regardless of unit of data analysis. The former addresses changes at the cluster level (either based on a cluster-level outcome or aggregated individual-level outcomes) and equally weights each cluster regardless of cluster size; the latter addresses changes for the population of individuals pooled across clusters and will upweight larger clusters.
@@
    (Mathematical difference) The cluster-ATE and individual-ATE estimands will be different in magnitude when the cluster size is informative, i.e., when the cluster size is marginally associated with the outcome or the treatment contrast.4, 6 If the cluster size is non-informative and individual outcomes are marginally identically distributed within a cluster,$\mu_{C}(a)=\mu_{I}(a)$ and $\Delta_{C}=\Delta_{I}$ and these two estimands coincide.22
@@ 3.1 Constructing the estimator
In a CRT, it is impossible to fully observe $\{Y_{ij}(1),Y_{ij}(0)\}$ for each individual, so that the estimand $\mu_{\omega}(a)$ is not directly estimable. We describe two assumptions that are plausible in CRTs. First, the observed outcome $Y_{ij}$, under the cluster-level Stable Unit Treatment Value Assumption, can be expressed as $Y_{ij}=A_{i}Y_{ij}(1)+(1-A_{i})Y_{ij}(0)$. Second, we assume cluster-level randomization such that $A_{i}\perp\{\bm{Y}_{i}(1),\bm{Y}_{i}(0)\}|\{\bm{X}_{i},\bm{H}_{i},N_{i}\}$, where, for cluster $i$, $\bm{Y}_{i}(a)=\left(Y_{i1}(a),\ldots,Y_{iN_{i}}(a)\right)^{T}$, $\bm{X}_{i}=\left(\bm{X}_{i1},\ldots,\bm{X}_{iN_{i}}\right)^{T}$, $\bm{H}_{i}$ is the collection of cluster-level covariates defined earlier, and $N_{i}$ is the cluster size. The randomization assumption holds by design, and includes as special cases (R1) simple randomization (where cluster-level randomization is determined by flipping a coin), (R2) stratified randomization23 (where one or more discrete cluster-level factor is considered to stratify clusters prior to randomization); (R3) pair-matching9 (where randomization is done within pairs of clustered formed based on similar cluster-level attributes) and (R4) covariate-constrained randomization24, 25 (where randomization is conducted within the subspace where sufficient balance is achieved according to baseline covariates). Under randomization schemes (R2)-(R4), the randomization of $A_{i}$ is conducted within levels of baseline covariates and the randomization probability may or may not vary according to levels of baseline covariates.
@@
First, $\widehat{\mu}_{\omega}(a)$ is a weighted average of the cluster-specific contributions, with the weight proportional to $\omega_{i}$. Each cluster-specific contribution takes the form of a regression prediction $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$, plus a weighted cluster-level residual. This is different from a purely model-based standardization estimator, $\sum_{i=1}^{m}\frac{\omega_{i}}{\omega_{+}}\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$, which is consistent only when the regression model is correctly specified.26 Due to the presence of a weighted residual term in (4), $\widehat{\mu}_{W}(a)$ is consistent and asymptotically normal for estimating $\Delta_{\omega}$, regardless of whether the outcome regression model for the cluster mean outcome $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ matches the true data generating process (Web Appendix A). This property then offers sufficient flexibility in model specification for $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$. In other words, one can use (4) as a unified solution to standardize the output from any fitted regression model for estimand-aligned inference in CRTs. By standardizing the regression model output, we mean taking the predicted outcome values from a fitted model and averaging in a manner aligned with the target estimand; Section 3.3 offers concrete examples.
Second, we show in the Web Appendix A that $\widehat{\mu}_{\omega}(a)$ is the solution to the unbiased estimating equations motivated by the efficient influence function27, 28 of $\mu_{\omega}(a)$. This implies that, if the outcome model for $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ matches the true data generating process, then $\widehat{\mu}_{\omega}(a)$ will be fully efficient as the large-sample variance of $\widehat{\mu}_{\omega}(a)$ achieves the semiparametric efficiency lower bound of $\mu_{\omega}(a)$. Such an estimator can substantially improve the efficiency over the unadjusted two-sample weighted difference-in-means estimator:
@@ 3.2 Specifying the randomization probability
Li et al.24, 25 demonstrated that covariate-constrained randomization is an effective randomization method to improve baseline balance in CRTs. As a recap, covariate-constrained randomization involves the following steps. First, one specifies key cluster-level attributes that need to be balanced by randomization. Second, one characterizes a balance metric that defines whether a proposed randomization scheme is acceptable in terms of baseline balance between groups. Third, one either enumerates (when the number of clusters is relatively limited) all or simulates (when there is a large number of clusters) a large number of potential randomization schemes based on the available clusters; in the case of simulated randomization schemes, the duplicate schemes are removed. Finally, one determines the threshold for acceptable randomization schemes, and choose a constrained randomization space as a subset of schemes where sufficient covariate balance is achieved according to the threshold. To specify $\pi(\bm{X}_{i},\bm{H}_{i},N_{i})$ under covariate-constrained randomization, we first let $R$ represent the number of all possible randomization schemes or the number of simulated unique schemes under constrained randomization. We let $m_{1}$, $m_{0}$ be the number of clusters assigned to treatment and control arms, respectively, and $m=m_{1}+m_{0}$. We then specify the $R\times m$ matrix $\bm{T}$, so that each row $[\bm{T}]_{r\cdot}=(t_{r1},t_{r2},\ldots,t_{rn})$ indicates a unique randomization scheme, where the $(r,i)$th element $t_{ri}$ is $1$ if cluster $i$ is randomized to intervention according to the $r$th scheme and 00 otherwise. Based on this constrained randomization matrix, one can approximate the vector of the randomization probabilities with $\left\{\pi(\bm{X}_{1},\bm{H}_{1},N_{1}),\pi(\bm{X}_{2},\bm{H}_{2},N_{2}),\ldots,\pi(\bm{X}_{m},\bm{H}_{m},N_{m})\right\}=\bm{1}_{R}^{T}\bm{T}/R$. In the special case that $m/2$ clusters are randomized to each condition and the balance metric in a symmetric quadratic form (such as the Mahalanobis distance), it is often sufficient to use $\pi(\bm{X}_{i},\bm{H}_{i},N_{i})=0.5$.
@@ 3.3 Specifying the outcome regression model
In principle, we can compute $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ using any parametric or semiparametric regression models designed for clustered data. We give several familiar examples on how this term might be obtained. We present models by adjusting for all covariates included in $\{\bm{X}_{i},\bm{H}_{i},N_{i}\}$; this is not necessary, as the consistency of our standardization estimator does not have stringent requirement on any of these outcome regression models. For example, one could omit the adjustment for cluster size or any other baseline covariates without affecting consistency. In practice, we carry the same recommendations made for individually randomized trials that all baseline covariates that adjusted for in the analysis should be pre-specified and include those that are prognostic of the outcome.29, 30 Finally, in all of the examples presented below, we consider only the main effects as linear terms (akin to the ANCOVA1 specification in Wang et al.22), but more flexible regression models including higher-order terms and interaction terms can be easily specified with minor modification.
@@
where $\overline{\bm{X}}_{i}=N_{i}^{-1}\sum_{j=1}^{N_{i}}\bm{X}_{ij}$ is the cluster-level mean covariate, and a working assumption is that the error $\epsilon_{i}$ follows a mean-zero normal distribution with a constant variance. This working assumption need not be correct (as $\overline{Y}_{i}$ is often heteroscedastic with variable cluster sizes) and is primarily to facilitate the ordinary least squares estimation. Notably, the cluster-level linear regression can be easily operationalized even if $Y_{ij}$ is binary or count, as the error assumption is only a working device. More generally, this approach has been known as the two-stage approach in the analysis of CRTs.31, 32 From (8), it is immediate that
@@
As a dominating approach in analyzing CRTs, the linear mixed model fitted with individual-level observations22 can be represented by
@@
where $\bm{\gamma}_{2}$ and $\bm{\gamma}_{3}$ separately account for the within-cluster and between-cluster effects of the individual-level covariates. In the analysis of clustered data, this model has been referred to as the contextual effect model; see, for example, Raudenbush33 and Tong et al.34 Setting $\bm{\gamma}_{2}=\bm{\gamma}_{3}$ eliminates the contextual effect in the model formulation. In (10), $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2})$ is the cluster-level random intercept and $\epsilon_{ij}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ is the random error. The outcome ICC given treatment and covariates, implied by the linear mixed model, is $\rho=\sigma_{b}^{2}/(\sigma_{b}^{2}+\sigma_{\epsilon}^{2})$. This quantity is of great interest in CRTs and plays a critical role in determining the sample size and power.2 Per the CONSORT extension of CRTs,8 reporting the ICC values in completed trials may facilitate the planning of future studies with a similar endpoint and has been a recommended practice. As in Example 3.1, the linear mixed model can be fit to binary and count data because the distributional assumptions for the random effect and the error term are only working assumptions; we refer to Wang et al.22 for a rigorous treatment of properties of the linear mixed models for the analysis of CRTs without informative cluster size. Under model (10), we have
@@
[…]
where $\phi(.)$ denotes the distribution of the random effect $c$. The last expression (12) is essentially a logistic-normal integral that can either be numerically approximated with arbitrary accuracy (see, for example, Crouch and Spiegelman35), or evaluated using Monte Carlo approximation. Since typically only a single random intercept is considered, a close approximation of (12) is given by the simpler expression ($\Pi$ is the mathematical constant)36
@@
That is, in log linear mixed models, marginalization over the random intercept only implies an overall intercept shift.37 The coefficient estimator for each generalized linear mixed model can be obtained from maximizing the approximate likelihood, through either the approach of penalized quasi-likelihood, Laplace approximation or adaptive Gaussian-quadrature;25 all three approaches have been implemented in the glmer function in the lme4 R package.
@@
Unlike the conditional models, marginal models obviate the need to condition on the latent random intercept. Instead, a generalized linear model is specified for the marginal mean (marginal in the sense that it is not a cluster-specific model with latent random effects), and a separate working correlation model is specified to account for the within-cluster associations between the outcomes.38 A general form of the marginal mean model is given by
@@
which does not involve any integration regardless of the link function. The estimation routine for the regression coefficients is based on solving the GEE, and has been extensively studied for clustered data; see, Preisser et al.39 for a detailed discussion on the merits of GEE for the design and analysis of parallel-arm CRTs. The implementation of GEE requires specifying the within-cluster correlation structure. Denoting $\bm{C}_{i}=\text{corr}(Y_{i1},\ldots,Y_{iN_{i}})$ as the working correlation structure for cluster $i$, the GEE regression estimator in CRTs can be based on either the independence working structure ($\bm{C}_{i}=\bm{I}_{N_{i}}$), exchangeable working structure ($\bm{C}_{i}=(1-\rho)\bm{I}_{N_{i}}+\rho\bm{1}_{N_{i}}\bm{1}^{T}_{N_{i}}$ with $\rho$ being the ICC), or arm-specific exchangeable working structure ($\bm{C}_{i}=\{A_{i}(1-\rho_{1})+(1-A_{i})(1-\rho_{0})\})\bm{I}_{N_{i}}+\{A_{i}\rho_{1}+(1-A_{i})\rho_{0}\}\bm{1}_{N_{i}}\bm{1}^{T}_{N_{i}}$ with $\rho_{a}$ being the arm-specific ICC).40 For a correctly specified marginal mean model, specifying an approximately correct working correlation structure generally leads to improved efficiency in estimating the regression coefficients compared to the independence working structure.41 To improve the standard GEE implementation, Preisser et al.42 provided matrix-adjusted estimating equations (MAEE) to correct for the negative bias in the correlation estimates in the presence of a small number of clusters. Finally, more flexible specifications of the working correlation structure can be considered following the quadratic inference functions approach.43, 44 As a further example, the marginalized multilevel models exploit the strengths of the generalized linear mixed model (in Example 3.3) to specify flexible within-cluster correlations but still enable the estimation of the marginal mean model parameters in (14).45 The tools for implementing marginal models are also developed. For instance, the general GEE routine can be implemented in the geeglm function in the geepack R package, whereas the MAEE routine can be found in the geeCRT R package. R code for implementing the quadratic inference function is available from the Appendix of Westgate.43

### v2

@@ 1 Introduction
Several prior studies have investigated estimators that target the cluster-ATE or individual-ATE as two marginal estimands in CRTs. Without baseline covariate adjustment, Table 3 in Kahan et al.4 provided a recipe for consistent estimators using either the individual-level data or cluster-summary data; they recommended the (weighted) independence generalized estimating equations (GEE) when the cluster size is informative. Wang et al.11 pointed out that the probability limit of the treatment coefficient estimator from GEE with an exchangeable working correlation structure is neither the cluster-ATE nor the individual-ATE. Similar findings also apply to parallel-arm CRTs with a baseline period.lee2025should With covariate adjustment, Su and Ding 12 established the asymptotic theory for a class of linearly-adjusted analysis of covariance estimators for estimating the weighted average treatment effects in CRTs, assuming working independence. Balzer et al.13, 14 and Benitez et al.10 studied targeted maximum likelihood estimators with machine learning methods for covariate adjustment in CRTs, and extended their methods to address missing outcomes;15 A further extension allowing for covariates missing not at random is given in Wang et al.16 Wang et al.17 developed efficient covariate-adjusted estimators for estimating both the cluster-ATE and individual-ATE, allowing for informative within-cluster subsampling. Despite these recent developments, the most common practice for analyzing CRTs relies on fitting regression models suitable for clustered data, including the (generalized) linear mixed models and GEE. For example, Fiero et al.18 reviewed 86 CRTs published between August 2013 and July 2014 and found that 52% used mixed models and 16% used GEE in their primary analyses. In a review of 100 CRTs from the United Kingdom’s National Institute for Health Research online Journals Library (from January 1997 to July 2021), 80% of the studies adopted a linear mixed or generalized linear mixed model for data analysis.19 For practitioners, guidance has been limited on how to draw robust inference about the cluster-ATE and the individual-ATE using common regression models for clustered data, irrespective of choices regarding covariate adjustment, working correlation and random-effects specification.
@@ 2 Formalizing the treatment effect estimands
We consider a standard parallel-arm CRT, where individuals are nested within clusters. We assume that data observations are collected through $m$ clusters, each of which contains $N_{i}$ individuals that we include in a study. We assume the cluster size $N_{i}$ is the study population of interest in each cluster and no further within-cluster subsampling is considered; methods that address the generalization to potentially unobserved cluster population can be found in Wang et al.17 and may require the additional knowledge of the source population size in each cluster. We let $A_{i}\in\{0,1\}$ be the randomized cluster-level treatment indicator, with $A_{i}=1$ indicating the assignment to the treatment condition and $A_{i}=0$ to usual care. Under the potential outcomes framework, we define $\{Y_{ij}(1),Y_{ij}(0)\}$ as a pair of potential outcomes for each individual $j\in\{1,\dots,N_{i}\}$ under the treatment and usual care conditions, respectively. Writing $f(a,b)$ as a pre-specified contrast function, we define a general class of weighted average treatment effect in CRTs as
@@
Although estimand (3) does not appear to be an individual-average quantity from a first look, its finite-population analogue can be defined as $\sum_{i=1}^{m}\sum_{j=1}^{N_{i}}Y_{ij}(a)/\sum_{i=1}^{m}N_{i}$ (by replacing the population-level expectation with an empirical one)12 and clearly reveals its feature. The individual-ATE addresses the expected change in outcome associated with treatment across the population of individuals across clusters, and can be relevant when the intervention is targeting either the cluster or the individual (despite cluster randomization).
@@
    (Conceptual difference) The cluster-ATE and the individual-ATE estimands focus on different units of inference,20 regardless of unit of data analysis. The former addresses changes at the cluster level (either based on a cluster-level outcome or aggregated individual-level outcomes) and equally weights each cluster regardless of cluster size; the latter addresses changes for the population of individuals pooled across clusters and will upweight larger clusters.
@@
[…]
@@ 3.1 Constructing the estimator
In a CRT, it is impossible to fully observe $\{Y_{ij}(1),Y_{ij}(0)\}$ for each individual, so that the estimand $\mu_{\omega}(a)$ is not directly estimable. We describe two assumptions that are plausible in CRTs. First, the observed outcome $Y_{ij}$, under the cluster-level Stable Unit Treatment Value Assumption, can be expressed as $Y_{ij}=A_{i}Y_{ij}(1)+(1-A_{i})Y_{ij}(0)$. Second, we assume cluster-level randomization such that $A_{i}\perp\{\bm{Y}_{i}(1),\bm{Y}_{i}(0)\}|\{\bm{X}_{i},\bm{H}_{i},N_{i}\}$, where, for cluster $i$, $\bm{Y}_{i}(a)=\left(Y_{i1}(a),\ldots,Y_{iN_{i}}(a)\right)^{T}$, $\bm{X}_{i}=\left(\bm{X}_{i1},\ldots,\bm{X}_{iN_{i}}\right)^{T}$, $\bm{H}_{i}$ is the collection of cluster-level covariates defined earlier, and $N_{i}$ is the cluster size. The randomization assumption holds by design, and includes as special cases (R1) simple randomization (where cluster-level randomization is determined by flipping a coin), (R2) stratified randomization22 (where one or more discrete cluster-level factor is considered to stratify clusters prior to randomization); (R3) pair-matching9 (where randomization is done within pairs of clustered formed based on similar cluster-level attributes) and (R4) covariate-constrained randomization23, 24 (where randomization is conducted within the subspace where sufficient balance is achieved according to baseline covariates). Under randomization schemes (R2)-(R4), the randomization of $A_{i}$ is conducted within levels of baseline covariates and the randomization probability may or may not vary according to levels of baseline covariates.
@@
First, $\widehat{\mu}_{\omega}(a)$ is a weighted average of the cluster-specific contributions, with the weight proportional to $\omega_{i}$. Each cluster-specific contribution takes the form of a regression prediction $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$, plus a weighted cluster-level residual. This is different from a purely model-based standardization estimator, $\sum_{i=1}^{m}\frac{\omega_{i}}{\omega_{+}}\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$, which is consistent only when the regression model is correctly specified.25 Due to the presence of a weighted residual term in (4), $\widehat{\mu}_{W}(a)$ is consistent and asymptotically normal for estimating $\Delta_{\omega}$, regardless of whether the outcome regression model for the cluster mean outcome $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ matches the true data generating process (Web Appendix A). This property then offers sufficient flexibility in model specification for $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$. In other words, one can use (4) as a unified solution to standardize the output from any fitted regression model for estimand-aligned inference in CRTs. By standardizing the regression model output, we mean taking the predicted outcome values from a fitted model and averaging in a manner aligned with the target estimand; Section 3.3 offers concrete examples.
Second, we show in the Web Appendix A that $\widehat{\mu}_{\omega}(a)$ is the solution to the unbiased estimating equations motivated by the efficient influence function26, 27 of $\mu_{\omega}(a)$. This implies that, if the outcome model for $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ matches the true data generating process, then $\widehat{\mu}_{\omega}(a)$ will be fully efficient as the large-sample variance of $\widehat{\mu}_{\omega}(a)$ achieves the semiparametric efficiency lower bound of $\mu_{\omega}(a)$. Such an estimator can substantially improve the efficiency over the unadjusted two-sample weighted difference-in-means estimator:
@@ 3.2 Specifying the randomization probability
Li et al.23, 24 demonstrated that covariate-constrained randomization is an effective randomization method to improve baseline balance in CRTs. As a recap, covariate-constrained randomization involves the following steps. First, one specifies key cluster-level attributes that need to be balanced by randomization. Second, one characterizes a balance metric that defines whether a proposed randomization scheme is acceptable in terms of baseline balance between groups. Third, one either enumerates (when the number of clusters is relatively limited) all or simulates (when there is a large number of clusters) a large number of potential randomization schemes based on the available clusters; in the case of simulated randomization schemes, the duplicate schemes are removed. Finally, one determines the threshold for acceptable randomization schemes, and choose a constrained randomization space as a subset of schemes where sufficient covariate balance is achieved according to the threshold. To specify $\pi(\bm{X}_{i},\bm{H}_{i},N_{i})$ under covariate-constrained randomization, we first let $R$ represent the number of all possible randomization schemes or the number of simulated unique schemes under constrained randomization. We let $m_{1}$, $m_{0}$ be the number of clusters assigned to treatment and control arms, respectively, and $m=m_{1}+m_{0}$. We then specify the $R\times m$ matrix $\bm{T}$, so that each row $\bm{T}]_{r\cdot}=(t_{r1},t_{r2},\ldots,t_{rn})$ indicates a unique randomization scheme, where the $(r,i)$th element $t_{ri}$ is $1$ if cluster $i$ is randomized to intervention according to the $r$th scheme and 0 otherwise. Based on this constrained randomization matrix, one can approximate the vector of the randomization probabilities with $\left\{\pi(\bm{X}_{1},\bm{H}_{1},N_{1}),\pi(\bm{X}_{2},\bm{H}_{2},N_{2}),\ldots,\pi(\bm{X}_{m},\bm{H}_{m},N_{m})\right\}=\bm{1}_{R}^{T}\bm{T}/R$. In the special case that $m/2$ clusters are randomized to each condition and the balance metric in a symmetric quadratic form (such as the Mahalanobis distance), it is often sufficient to use $\pi(\bm{X}_{i},\bm{H}_{i},N_{i})=0.5$.
@@ 3.3 Specifying the outcome regression model
In principle, we can compute $\widehat{E}(\overline{Y}_{i}|A_{i}=a,\bm{X}_{i},\bm{H}_{i},N_{i})$ using any parametric or semiparametric regression models designed for clustered data. We give several familiar examples on how this term might be obtained. We present models by adjusting for all covariates included in $\{\bm{X}_{i},\bm{H}_{i},N_{i}\}$; this is not necessary, as the consistency of our standardization estimator does not have stringent requirement on any of these outcome regression models. For example, one could omit the adjustment for cluster size or any other baseline covariates without affecting consistency. In practice, we carry the same recommendations made for individually randomized trials that all baseline covariates that adjusted for in the analysis should be pre-specified and include those that are prognostic of the outcome.28, 29 Finally, in all of the examples presented below, we consider only the main effects as linear terms (akin to the ANCOVA1 specification in Wang et al.21), but more flexible regression models including higher-order terms and interaction terms can be easily specified with minor modification.
@@
where $\overline{\bm{X}}_{i}=N_{i}^{-1}\sum_{j=1}^{N_{i}}\bm{X}_{ij}$ is the cluster-level mean covariate, and a working assumption is that the error $\epsilon_{i}$ follows a mean-zero normal distribution with a constant variance. This working assumption need not be correct (as $\overline{Y}_{i}$ is often heteroscedastic with variable cluster sizes) and is primarily to facilitate the ordinary least squares estimation. Notably, the cluster-level linear regression can be easily operationalized even if $Y_{ij}$ is binary or count, as the error assumption is only a working device. More generally, this approach has been known as the two-stage approach in the analysis of CRTs.30, 31 From (8), it is immediate that
@@
As a dominating approach in analyzing CRTs, the linear mixed model fitted with individual-level observations21 can be represented by
@@
where $\bm{\gamma}_{2}$ and $\bm{\gamma}_{3}$ separately account for the within-cluster and between-cluster effects of the individual-level covariates. In the analysis of clustered data, this model has been referred to as the contextual effect model; see, for example, Raudenbush32 and Tong et al.33 Setting $\bm{\gamma}_{2}=\bm{\gamma}_{3}$ eliminates the contextual effect in the model formulation. In (10), $b_{i}\sim\mathcal{N}(0,\sigma_{b}^{2})$ is the cluster-level random intercept and $\epsilon_{ij}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ is the random error. The outcome ICC given treatment and covariates, implied by the linear mixed model, is $\rho=\sigma_{b}^{2}/(\sigma_{b}^{2}+\sigma_{\epsilon}^{2})$. This quantity is of great interest in CRTs and plays a critical role in determining the sample size and power.2 Per the CONSORT extension of CRTs,8 reporting the ICC values in completed trials may facilitate the planning of future studies with a similar endpoint and has been a recommended practice. As in Example 3.1, the linear mixed model can be fit to binary and count data because the distributional assumptions for the random effect and the error term are only working assumptions; we refer to Wang et al.21 for a rigorous treatment of properties of the linear mixed models for the analysis of CRTs without informative cluster size. Under model (10), we have
@@
$=\frac{1}{N_{i}}\sum_{i=1}^{N_{i}}\int_{0}^{\infty}\text{expit}\left\{\widehat{\alpha}_{0}+\widehat{\alpha}_{1}A_{i}+\widehat{\bm{\alpha}}_{2}^{T}(\bm{X}_{ij}-\overline{\bm{X}}_{i})+\widehat{\bm{\alpha}}_{3}^{T}\overline{\bm{X}}_{i}+\widehat{\bm{\alpha}}_{4}^{T}{\bm{H}}_{i}+\widehat{\alpha}_{5}N_{i}+c\right\}\phi(c|0,\widehat{\sigma}_{c}^{2})dc,$ (12)
where $\phi(.)$ denotes the distribution of the random effect $c$. The last expression (12) is essentially a logistic-normal integral that can either be numerically approximated with arbitrary accuracy (see, for example, Crouch and Spiegelman34), or evaluated using Monte Carlo approximation. Since typically only a single random intercept is considered, a close approximation of (12) is given by the simpler expression ($\Pi$ is the mathematical constant)35
@@
That is, in log linear mixed models, marginalization over the random intercept only implies an overall intercept shift.36 The coefficient estimator for each generalized linear mixed model can be obtained from maximizing the approximate likelihood, through either the approach of penalized quasi-likelihood, Laplace approximation or adaptive Gaussian-quadrature;24 all three approaches have been implemented in the glmer function in the lme4 R package.
@@
Unlike the conditional models, marginal models obviate the need to condition on the latent random intercept. Instead, a generalized linear model is specified for the marginal mean (marginal in the sense that it is not a cluster-specific model with latent random effects), and a separate working correlation model is specified to account for the within-cluster associations between the outcomes.37 A general form of the marginal mean model is given by
@@
which does not involve any integration regardless of the link function. The estimation routine for the regression coefficients is based on solving the GEE, and has been extensively studied for clustered data; see, Preisser et al.38 for a detailed discussion on the merits of GEE for the design and analysis of parallel-arm CRTs. The implementation of GEE requires specifying the within-cluster correlation structure. Denoting $\bm{C}_{i}=\text{corr}(Y_{i1},\ldots,Y_{iN_{i}})$ as the working correlation structure for cluster $i$, the GEE regression estimator in CRTs can be based on either the independence working structure ($\bm{C}_{i}=\bm{I}_{N_{i}}$), exchangeable working structure ($\bm{C}_{i}=(1-\rho)\bm{I}_{N_{i}}+\rho\bm{1}_{N_{i}}\bm{1}^{T}_{N_{i}}$ with $\rho$ being the ICC), or arm-specific exchangeable working structure ($\bm{C}_{i}=\{A_{i}(1-\rho_{1})+(1-A_{i})(1-\rho_{0})\})\bm{I}_{N_{i}}+\{A_{i}\rho_{1}+(1-A_{i})\rho_{0}\}\bm{1}_{N_{i}}\bm{1}^{T}_{N_{i}}$ with $\rho_{a}$ being the arm-specific ICC).39 For a correctly specified marginal mean model, specifying an approximately correct working correlation structure generally leads to improved efficiency in estimating the regression coefficients compared to the independence working structure.40 To improve the standard GEE implementation, Preisser et al.41 provided matrix-adjusted estimating equations (MAEE) to correct for the negative bias in the correlation estimates in the presence of a small number of clusters. Finally, more flexible specifications of the working correlation structure can be considered following the quadratic inference functions approach.42, 43 As a further example, the marginalized multilevel models exploit the strengths of the generalized linear mixed model (in Example 3.3) to specify flexible within-cluster correlations but still enable the estimation of the marginal mean model parameters in (14).44 The tools for implementing marginal models are also developed. For instance, the general GEE routine can be implemented in the geeglm function in the geepack R package, whereas the MAEE routine can be found in the geeCRT R package. R code for implementing the quadratic inference function is available from the Appendix of Westgate.42